# Nonlinear subspace clustering by functional link neural networks

### v1

@@ 5.1 Experimental settings
1) Datasets: We consider four publicly available datasets: Extended Yale B wang2016constrained , USPS yang2019subspace , COIL20 and and ORL lu2018subspace . For these three datasets, we use PCA to perform dimensionality on the raw data, with the default dimension set to $\rm{nCluster}\times 6$. A brief description of each dataset is provided below:
@@
2) Compared methods: We compare FLNNSC and CCSC with several baselines, including SSC elhamifar2013sparse , LRR liu2012robust , LRSC vidal2014low , NSC zhu2018nonlinear , SMR hu2014smooth , KSSC patel2014kernel , LSR1 and LSR2 lu2012robust , Kernel Truncated Regression Representation (KTRR) zhen2020kernel , BDR-B and BDR-Z lu2018subspace .
@@ 5.2.1 Performance comparison
Across all experiments with different subjects, the proposed FLNNSC and CCSC algorithms consistently achieve higher clustering accuracy than all other compared subspace clustering methods. Furthermore, CCSC exhibits superior performance compared to FLNNSC, with an accuracy improvement of $1.8\%$ and $2.5\%$ for 10 subjects and 15 subjects, respectively, attributable to the combination parameter $\lambda$ for balancing the influence of linear and nonlinear representations. When considering 20 subjects, both CCSC and FLNNSC achieve the same clustering accuracy, indicating that the optimal performance of CCSC is attained when $\lambda=1$, thus demonstrating that CCSC reduces to the FLNNSC algorithm in this case.
Regarding other metrics, such as NMI, ARI and F1 score, our proposed algorithms consistently outperform other subspace clustering methods, except that the ARI and F1-score metrics are inferior to those of KTRR in the case of 20 subjects.
For the purpose of exploring the influence of the combination parameter $\lambda$ on CCSC, we plot the evolutionary curves of CCSC with respect to $\lambda$, while maintaining $\alpha$ and $\beta$ fixed at the optimal parameters of FLNNSC. The plots are shown in Figure 2. From Figure 2(a), we observe that CCSC exhibits relatively low clustering accuracy for small $\lambda$ values, where linear representations predominantly contribute to the results. As $\lambda$ increases, the performance improves, reaching its peak when $\lambda=1$, where the nonlinear representations are fully exploited without any influence from linear representations. Moreover, a similar observation can be made from Figure 2(b), except NMI experiencing a slight decline from $\lambda=0.9$ to $\lambda=1$.
@@ 5.2.3 Parameter sensitivity and convergence analysis
2) Convergence analysis: We conduct an experiment analysis to examine the convergence behavior of the proposed algorithm, utilizing the metric $\lVert\mathbf{Z}_{k}-\mathbf{Z}_{k-1}\rVert^{2}_{F}$, as shown in Figure 5. The obtained results demonstrate that the proposed FLNNSC algorithm exhibits rapid convergence across all three datasets, reaching a steady-state in a remarkably short period.
@@ 5.2.4 Execution time analysis
In order to validate the efficiency of our method, we choose two benchmark methods for comparison: NSC and BDR. We select NSC for comparison as our method is developed in response to the limitations inherent in this particular approach. On the other hand, BDR is selected as it is a well-established baseline method known for its impressive performance across a variety of datasets. Figure 6 shows the comparison results of execution time on different datasets. It is seen that as the number of clusters increases, our method consistently outperforms both NSC and BDR in terms of execution time, thereby validating the efficiency of our proposed method.
@@ 5.3 Discussion
In subspace clustering, when the data exhibits evident nonlinearity characteristics, the combination method, i.e., CCSC, shows comparable clustering performance to FLNNSC under the optimal parameter settings. However, if the data also exhibits significant linear characteristics that cannot be ignored, CCSC outperforms FLNNSC due to its ability to dynamically adjust the balance between linear and nonlinear representations. Moreover, the proposed FLNNSC algorithm, using a functional link neural network for nonlinear representation, has higher computational efficiency compared to traditional neural network-based subspace clustering methods. Remarkably, to enhance the capacity for capturing the data’s nonlinearity, one can employ a higher-order functional expansion for FLNN. This may lead to performance improvements, but this enhancement introduces a trade-off by increasing computational intricacy.
@@ 6 Conclusions
We proposed a novel nonlinear subspace clustering method named as FLNNSC in this paper. By utilizing a single-layer FLNN with remarkable nonlinearity approximation capability, FLNNSC demonstrates the ability to exploit the nonlinear features of data, resulting in enhanced clustering performance. Moreover, the proposed FLNNSC algorithm was regularized to enable the grouping effect. To exploit both linear and nonlinear characteristics, we further introduced the CCSC algorithm, which dynamically adjusts the balance between linear and nonlinear representations via the combination parameter. Experiments on benchmark datasets have shown that our methods outperform several advanced subspace clustering methods. Additionally, affinity graph results have highlighted the grouping effect of FLNNSC. We conducted a parameter sensitivity analysis, yielding valuable suggestions for parameter selection. Furthermore, we empirically demonstrated the rapid convergence behavior of FLNNSC.
@@ Acknowledgment
The work of Long Shi was partially supported by the National Natural Science Foundation of China under Grant 62201475. The work of Badong Chen was supported by the National Natural Science Foundation of China under Grants U21A20485 and 61976175. The work of Yu Zhao was supported by the Sichuan Science and Technology Program under Grant No. 2023NSFSC0032.

### v2

@@ 5.1 Experimental settings
1) Datasets: We consider four publicly available datasets: Extended Yale B [48], USPS [49], COIL20 and and ORL [9]. We use PCA to perform dimensionality on the raw datasets. For implementing PCA, we reduce the data dimensionality to a default dimensionality. Since the determination of the number of principal components is influenced by the number of cluster, we set the number of principal components (default dimensionality) to be $clusters\times 6$. A brief description of each dataset is provided below:
@@
2) Compared methods: We compare FLNNSC and CCSC with several baselines, including SSC [3], LRR [5], LRSC [50], NSC [39], SMR [22], KSSC [31], LSR1 and LSR2 [20], Kernel Truncated Regression Representation (KTRR) [34], BDR-B and BDR-Z [9].
@@ 5.2 Evaluation of clustering performance
In the following, we will conduct extensive experiments to validate our proposed model, including clustering performance comparison, affinity graph analysis, parameter sensitivity and convergence analysis, as well as execution time analysis.
@@ 5.2.1 Performance comparison
It is seen that two variants based on BDR, namely BDR-B and BDR-Z, outperform other existing subspace clustering methods in performance. This superior performance can be attributed to their direct enforcement of a block diagonal structure, which theoretically assures desirable clustering results. Meanwhile, the SMR method, which incorporates a regularization for the grouping effect, also demonstrates admirable performance. Furthermore, we notice that KSSC outperforms SSC, which verifies the efficacy of kernel mapping in effectively exploring the nonlinear structure among data.
Across all experiments with different subjects, the proposed FLNNSC and CCSC algorithms consistently achieve higher clustering accuracy than all other compared subspace clustering methods. This can be attributed to the fact that both of our proposed methods incorporate considerations for nonlinear transformation and the grouping effect during the modeling process. As a result, they theoretically ensure optimal clustering performance when dealing with practical datasets. Furthermore, CCSC exhibits superior performance compared to FLNNSC, with an accuracy improvement of $1.8\%$ and $2.5\%$ for 10 subjects and 15 subjects, respectively. This enhanced performance is attributable to the combination parameter $\lambda$, which effectively balances the influence of linear and nonlinear representations. When considering 20 subjects, both CCSC and FLNNSC achieve the same clustering accuracy, indicating that the optimal performance of CCSC is attained when $\lambda=1$, thus demonstrating that CCSC reduces to the FLNNSC algorithm in this case. Regarding other metrics, such as NMI, ARI and F1 score, our proposed algorithms consistently outperform other subspace clustering methods, except that the ARI and F1-score metrics are inferior to those of KTRR in the case of 20 subjects.
@@
For the purpose of exploring the influence of the combination parameter $\lambda$ on CCSC, we plot the evolutionary curves of CCSC with respect to $\lambda$, while maintaining $\alpha$ and $\beta$ fixed at the optimal parameters of FLNNSC. The plots are shown in Figure 2. From Figure 2(a), we observe that CCSC exhibits relatively low clustering accuracy for small $\lambda$ values, where linear representations predominantly contribute to the results. As $\lambda$ increases, the performance improves, reaching its peak when $\lambda=1$, where the nonlinear representations are fully exploited without any influence from linear representations. In this particular case where $\lambda=1$, CCSC actually reduces to FLNNSC. This implies that, on the USPS dataset, the optimal performance of CCSC aligns precisely with that of FLNNSC. This result further emphasizes the significance of investigating the nonlinear structure within data. Moreover, a similar observation can be made from Figure 2(b), except NMI experiencing a slight decline from $\lambda=0.9$ to $\lambda=1$. Overall, focusing more on uncovering the nonlinear characteristics of data is beneficial to subspace clustering.
@@ 5.2.3 Parameter sensitivity and convergence analysis
2) Convergence analysis: We conduct an experiment analysis to examine the convergence behavior of the proposed algorithm, utilizing the metric $\lVert\mathbf{Z}_{k}-\mathbf{Z}_{k-1}\rVert^{2}_{F}$, as shown in Figure 5. It is observed that FLNNSC achieves a steady-state within 10 iterations. Specifically, on the Extended Yale B and COIL20 datasets, it reaches this state in fewer than 5 iterations. The obtained results demonstrate that the proposed FLNNSC algorithm exhibits rapid convergence across all three datasets, reaching a steady-state in a remarkably short period. This rapid convergence feature not only significantly reduces time consumption but also caters to the real-time requirements in practical scenarios, especially when processing large-scale data or dealing with complex tasks.
@@ 5.2.4 Execution time analysis
In order to further validate the efficiency of our method, we choose two benchmark methods for comparison: NSC and BDR. We select NSC for comparison as our method is developed in response to the limitations inherent in this particular approach. On the other hand, BDR is selected as it is a well-established baseline method known for its impressive performance across a variety of datasets. Figure 6 shows the comparison results of execution time on different datasets. It is seen that as the number of clusters increases, our method consistently outperforms both NSC and BDR in terms of execution time, thereby validating the efficiency of our proposed method. It is expected that FLNNSC will exhibit promising performance in terms of execution efficiency in real-world applications.
@@
5.3 Friedman test
We consider Friedman test [51] to validate the superiority of the proposed model. Specifically, in terms of datasets, we select Extended Yale B with 10 subjects, USPS, COIL20, and ORL. Regarding methods, we choose SSC, LRR, LRSC, NSC, SMR, KSSC, LSR1, KTRR, BDR-B, and our FLNNSC. For each dataset, we sort various models in ascending order based on their accuracy and use the index of the sorted values as the rank for the corresponding model, and the rank matrix is shown in Table 7.
Table 7: Rank matrix of different methods on four datasets.
  Datasets SSC LRR LRSC NSC SMR KSSC LSR1 KTRR BDR-B FLNNSC
  Extended Yale B 2 5 7 1 6 4 3 8 9 10
USPS 4 7 1 5 8 2 3 6 9 10
COIL20 5 6 2 4 3 7 1 9 8 10
ORL 1 2 3 7 5 4 6 8 9 10
After obtaining the rank matrix, we calculate the rank value for each method. Let $rank_{i,j}$ be the rank of the $i$-th method on the $j$-th dataset, then the rank of the $i$-th model is:
$r_{i}=\frac{1}{N}\sum^{N}_{j=1}rank_{i,j},$ (24)
where $N$ denotes the number of datasets. Assuming that rank of each method follows a normal distribution, the corresponding chi-square statistic is [52]:
$\tau_{\chi^{2}}=\frac{12N}{k(k+1)}\left(\sum^{k}_{i=1}r^{2}_{i}-\frac{k(k+1)^{2}}{4}\right),$ (25)
where $k$ is the number of methods. Finally, we can use the chi-square distribution to compute the $p$-value. By applied the friedman function in Matlab, we calculate a $p$-value of $0.0039$. As this value is lower than the significance level of $0.05$, we reject the null hypothesis. This suggests that there is a significant difference in the performance of the various methods. To visually represent these differences, we have plotted a bar chart of mean rank, as shown in Fig. 7. It is evident from the chart that KTRR, BDR-B, and FLNNSC significantly outperform the other methods in terms of mean rank (calculated using equation (24)). Notably, our method, FLNNSC, achieves a higher rank value than both KTRR and BDR-B.
Figure 7: Friedman test results.
@@ 5.4 Discussion
In subspace clustering, when the data exhibits evident nonlinearity characteristics, the combination method, i.e., CCSC, shows comparable clustering performance to FLNNSC under the optimal parameter settings. If one select CCSC, it is crucial to carefully tune an extra balance parameter $\lambda$. This is because the selection of $\lambda$ impacts the performance. Therefore, in such cases, we recommend FLNNSC. If the data also exhibits significant linear characteristics that cannot be ignored, CCSC outperforms FLNNSC due to its ability to dynamically adjust the balance between linear and nonlinear representations. In this case, we recommend CCSC, with the tuning of $\lambda$ over the range of (0.5, 1).
Moreover, the proposed FLNNSC algorithm, using a functional link neural network for nonlinear representation, has higher computational efficiency compared to traditional neural network-based subspace clustering methods. Remarkably, to enhance the capacity for capturing the data’s nonlinearity, one can employ a higher-order functional expansion for FLNN. This may lead to performance improvements, but this enhancement introduces a trade-off by increasing computational intricacy.
@@ 6 Conclusions
We proposed a novel nonlinear subspace clustering method named as FLNNSC in this paper. By utilizing a single-layer FLNN with remarkable nonlinearity approximation capability, FLNNSC demonstrates the ability to exploit the nonlinear features of data, resulting in enhanced clustering performance. Moreover, the proposed FLNNSC algorithm was regularized to enable the grouping effect. To exploit both linear and nonlinear characteristics, we further introduced the CCSC algorithm, which dynamically adjusts the balance between linear and nonlinear representations via the combination parameter. Experiments on benchmark datasets have shown that our methods outperform several advanced subspace clustering methods. Additionally, affinity graph results have highlighted the grouping effect of FLNNSC. We conducted a parameter sensitivity analysis, yielding valuable suggestions for parameter selection. Furthermore, we empirically demonstrated the rapid convergence behavior of FLNNSC. It should be noted in the CCSC method, the combination parameter $\lambda$, which balances linear and nonlinear representations, plays a crucial role. However, selecting an optimal $\lambda$ still imposes a challenge, as an ideal choice would depend on the underlying data features, which are typically unavailable in practice. In our future work, we will investigate a dynamic scheme to adaptively adjust $\lambda$.
@@ Acknowledgment
The work of Long Shi was partially supported by the National Natural Science Foundation of China under Grant 62201475 and Natural Science Foundation of Sichuan Province under Grant 2024NSFSC1436. The work of Badong Chen was supported by the National Natural Science Foundation of China under Grants U21A20485 and 61976175. The work of Yu Zhao was supported by the Sichuan Science and Technology Program under Grant No. 2023NSFSC0032.